Serving at a speed of 170 km/h, a tennis player hits the ball at a height of 2.5 m and an angle θ below the horizontal. The service line is 11.9 m from the net, which is 0.91 m high. What is the angle θ such that the ball just crosses the net? Will the ball land in the service box, whose out line is 6.40 m from the net?
1
Expert's answer
2021-10-28T08:53:51-0400

The horizontal distance to the net: "r=11.9 \text{ m}."

The horizontal component of velocity:


"v_x=v\cos\theta,\\\space\\
t=\frac r{v_x}=\frac r{v\cos\theta}."

Meanwhile, during the same time t, the ball must cover the height of


"h=2.5-0.91=1.19\text{ m}."

The vertical velocity at the net is


"u=v\sin\theta+gt."


The time is


"t=\frac{h}{v_\text{avg}}=\frac{2d}{2v\sin\theta+gt},\\\space\\
ht^2+2tv\sin\theta-2h=0,\\\space\\
\frac{hr^2}{v^2\cos^2\theta}+2v\tan\theta-2h=0,\\\space\\
\theta=0.05°."